An exit ramp of a highway is banked at an angle of 27 degrees


. The radius of ramp is 200m,


what is the maximum possible speed of an exiting car?



1
Expert's answer
2021-03-21T11:22:42-0400
"Nsin\theta=\dfrac{mv^2}{R}, (1)""Ncos\theta=mg. (2)"

Dividing equation (1) by (2), we get:


"tan\theta=\dfrac{v^2}{Rg},""v_{max}=\sqrt{Rgtan\theta},""v_{max}=\sqrt{200\ m\cdot9.8\ \dfrac{m}{s^2}\cdot tan27^{\circ}}=31.6\ \dfrac{m}{s}."